### v1

# Symplectic Bregman divergences

Abstract
We present a generalization of Bregman divergences in symplectic vector spaces called symplectic Bregman divergences. Symplectic Bregman divergences are derived from a symplectic generalization of the Fenchel-Young inequalities which rely on symplectic subdifferentials. The generic symplectic Fenchel-Young inequality is obtained using symplectic Fenchel transforms which are defined with respect to linear symplectic forms. Some potential appplications of symplectic divergences in geometric mechanics, information geometry, and learning dynamics in machine learning are discussed.
Keywords: duality product; inner product; linear symplectic form; symplectic matrix group; symplectic subdifferential; symplectic Fenchel transform; Moreau’s proximation; geometric mechanics.
1 Introduction
Symplectic geometry [9] was historically pioneered (locally and implicitly) by Lagrange around 1808 in his treatise on analytic mechanics [18, 19] where the motions and dynamics of a finite set of $m$ point mass particles in a time interval $T$ are analyzed in the phase space $\{(q_{1},p_{1},\ldots,q_{m},p_{m})(t){\ :\ }t\in T\}\subset\mathbb{R}^{2n}$, where $q_{i}(t)\in\mathbb{R}^{n}$’s denote the point locations at time $t$ and $p_{i}(t)\in\mathbb{R}^{n}$’s encode the momentum, i.e., $p_{i}(t)=m_{i}\dot{q}_{i}$ with $\dot{q}_{i}=\frac{d}{{\mathrm{d}t}}q_{i}(t)$. See Figure 1.
The Hamiltonian coupled equations governing the system motion are written in the phase space as
$\frac{{\mathrm{d}q}^{i}}{{\mathrm{d}t}}$ $=$ $\frac{\partial{H}}{\partial p_{i}},$ (1)
$\frac{{\mathrm{d}p}_{i}}{{\mathrm{d}t}}$ $=$ $-\frac{\partial H}{\partial q^{i}},$ (2)
where $H(q,p,t)$ is the Hamiltonian describing the system [12]. Symplectic geometry can be thought as the first discovered non-Euclidean geometric structure since hyperbolic geometry is usually considered to be first studied by Lobachevsky and Bolyai around the 1820’s-1930’s.
Figure 1: The motion of a single point particle $q(t)$ with mass $m$ and momentum $p(t)=m\dot{q}(t)$ on a 1D line can be modeled as a curve $\{(q(t),p(t)){\ :\ }t\in T\}$ in the phase space $\mathbb{R}^{2}$.
The adverb “symplectic” stems from Greek: It means “braided together” to convey the interactions of point mass particle positions with their momenta. Its use in mathematics originated in the work of Hermann Weyl [32]. Another synonym adverb of symplectic is “complex” which has been used to describe braided numbers $z$ of =$\mathbb{C}=\{z=a+ib{\ :\ }(a,b)\in\mathbb{R}^{2}\}$. Complex has its etymological root in Latin. In differential geometry, symplectic structures are closely related to complex structures on vector spaces and smooth manifolds [9].
In physics, symplectic geometry is not only at the core of classical mechanics (i.e., conservative reversible mechanics) and quantum mechanics [31], but has also recently been used to model and study dynamics of systems exhibiting dissipative terms [6] which are irreversible. As a pure geometry, symplectic geometry can be studied on its own by mathematicians, and gave birth to the field of symplectic topology [3]. Thus symplectic geometry can be fruitfully applied to various areas beyond its original domain of geometric mechanics. For example, symplectic geometry has been considered in machine learning for accelerating numerical optimization methods based on symplectic integrators [16] and in physics-informed neural networks [7, 20] (PINNs).
In this paper, we define symplectic Bregman divergences which recover as special cases composite Bregman divergences [5]. A Bregman divergence induced by a strictly convex and differentiable (potential) function $F$ is defined in [5] (1967) by
$B_{F}(x_{1}:x_{2})=F(x_{1})-F(x_{2})-{\langle x_{1}-x_{2},\nabla F(x_{2})\rangle}.$
When $F$ is separable, i.e. ,$F(z)=F_{1}(x)+F_{2}(y)$ for $z=(x,y)$, we say that $B_{F}$ is composite, and we have
$B_{F_{1},F_{2}}(z:z^{\prime})=B_{F_{1}}(x:x^{\prime})+B_{F_{2}}(y:y^{\prime}),$
where $z=(x,y)$ and $z^{\prime}=(x^{\prime},y^{\prime})$.
This paper introduces and extends the work of Buliga and Saxcé [6] which focused on geometric irreversible mechanics. To contrast with [6], this expository paper is targeted to an audience familiar with Bregman divergences [5] in machine learning and information geometry [1].
The paper is organized as follows: In §2, we define symplectic vector spaces and explain the representation of linear symplectic forms using either natural pairings or symplectic matrices. We then define the symplectic Fenchel transform and symplectic Fenchel-Young inequality in §3. The definitions of symplectic Fenchel-Young divergences and symplectic Bregman divergences are reported in §4. In particular, we show how to recover composite Bregman divergences as special cases in §5. Finally, we recall the interplay of Bregman divergences and dually flat manifolds of information geometry in §6, and motivate the introduction of symplectic Bregman divergences for geometric mechanics (BEN principle of [6]) and learning dynamics in ML.
2 Linear symplectic forms and symplectic groups
We begin with two definitions:
Definition 1 (Dual system.)
Let $X$ and $Y$ be topological locally convex $m$-dimensional vector spaces [14] equipped with a dual product:
${\langle\cdot,\cdot\rangle}:X\times Y\rightarrow\mathbb{R},$
such that all continuous linear functionals on $X$ and $Y$ are expressed as ${\langle x,\cdot\rangle}$ and ${\langle\cdot,y\rangle}$, respectively. The triplet $(X,Y,{\langle\cdot,\cdot\rangle})$ forms a dual system.
Definition 2 (Symplectic vector space)
A symplectic vector space $(V,\omega)$ is a vector space equipped with a map [22] $\omega:Z=V\times V\rightarrow\mathbb{R}$ which is
  1. 1.
    bilinear: $\forall\alpha,\beta,\alpha^{\prime},\beta^{\prime}\in\mathbb{R},\forall z_{1},z_{2}\in Z$, we have
    $\omega(\alpha z_{1}+\alpha^{\prime}z_{1}^{\prime},z_{2},\beta z_{2}+\beta^{\prime}z_{2}^{\prime})=\alpha\beta\,\omega(z_{1},z_{2})+\alpha\beta^{\prime}\,\omega(z_{1},z_{2}^{\prime})+\alpha^{\prime}\beta\,\omega(z_{1}^{\prime},z_{2})+\alpha^{\prime}\beta^{\prime}\,\omega(z_{1},z_{2}^{\prime}),$
  2. 2.
    skew-symmetric (or alternating): $\omega(z_{2},z_{1})=-\omega(z_{1},z_{2})$, and
  3. 3.
    non-degenerate: prescribed $\omega(z,z_{0})=0$ for all $z\in Z$ implies $z_{0}=0$.
Notice that skew-symmetry implies that $\omega(z,z)=0$ for all $z\in Z$ since $\omega(z,z)=-\omega(z,z)$ and hence $02\omega(z,z)=$ 0. The map $\omega$ is called a linear symplectic form.
Let us report several examples of linear symplectic forms:
  • Let $X=V$ as a finite $n$-dimensional vector space with the dual space of linear functionals (covectors $l$) $Y=V^{*}$. The natural pairing ${((x,l))}=l(x)=\sum_{i}x^{i}l_{i}$ of a vector $x\in V$ with a covector $l\in V^{*}$ is an example of dual product. We use the superscript index for indicating components of contravariant vectors and subscript index for specifying components of covariant vectors. We define the symplectic form induced by the natural pairing as follows:
    $\omega(z_{1},z_{2})$ $=$ ${((x_{1},y_{2}))}-{((x_{2},y_{1}))},$ (3)
    $=$ $l_{2}(x_{1})-l_{1}(x_{2}),$ (4)
    where $z_{1}=(x_{1},l_{1})$ and $z_{2}=(x_{2},l_{2})$ belong to $Z=V\oplus V^{*}$.
  • Consider $(X,{\langle\cdot,\cdot\rangle})$ an inner product space of dimension $n$. The product space $Z=X\oplus X$ of even dimension $n$ can be equipped with the following map $\omega:Z\times Z\rightarrow\mathbb{R}$ induced by the inner product:
    $\omega(z_{1},z_{2})={\langle x_{1},y_{2}\rangle}-{\langle x_{2},y_{1}\rangle},$ (5)
    where $z_{1}=(x_{1},y_{1})\in Z$ and $z_{2}=(x_{2},y_{2})\in Z$.
    Figure 2: Interpreting a 2D symplectic form $\omega(z_{1},z_{2})$ as the signed area of a parallelogram with first oriented edge $z_{1}$ (grey). A pair of vectors defines two possible orientations of the parallelogram: The orientation compatible with $z_{1}$ and the reverse orientation compatible with $z_{2}$.
    For example, let $X=\mathbb{R}$ and ${\langle x_{1},x_{2}\rangle}=x_{1}x_{2}$. Then $\omega(z_{1},z_{2})={x_{1}}{y_{2}}-{x_{2}}{y_{1}}$. The symplectic form can be interpreted as the determinant of the matrix $M={\left[\begin{array}[]{ll}x_{1}&x_{2}\cr y_{2}&y_{1}\end{array}\right]}$ which corresponds geometrically to the signed orientation of the parallelogram defined by the vectors $z_{1}=(x_{1},y_{1})$ and $z_{2}=(x_{2},y_{2})$. See Figure 2. This example indicates the link with integration of 2D manifolds equipped with fields of symplectic forms smoothly varying called differential $2$-forms [12].
    We can express the inner product as ${\langle x,y\rangle}=x^{\top}Qy$ for a symmetric positive-definite matrix $Q\in\mathbb{R}^{n\times n}$. Let $Q=L^{\top}L$ be the Cholesky decomposition of $Q$. Then we have ${\langle x,y\rangle}=(L^{\top}x)^{\top}I(L^{\top}y)={\langle L^{\top}x,L^{\top}y\rangle}_{0}$, where $I$ is the $n\times n$ identity matrix and ${\langle x,y\rangle}_{0}=x^{\top}y$ is the Euclidean inner product. Thus the form $\omega_{0}$ induced by ${\langle\cdot,\cdot\rangle}_{0}$ can be expressed using linear algebra as
    $\omega_{0}(z_{1},z_{2})=z_{1}^{\top}{\left[\begin{array}[]{ll}0&I\cr-I&0\end{array}\right]}z_{2}=z_{1}^{\top}\Omega_{0}z_{2},$
    where $\Omega_{0}\in\mathbb{R}^{2n\times 2n}$ is a skew-symmetric matrix: $\Omega_{0}^{\top}=-\Omega_{0}$. More generally, we may consider matrix skew-symmetric matrix $\Omega={\left[\begin{array}[]{ll}0&L^{\top}\cr-L^{\top}&0\end{array}\right]}$ to define the symplectic form $\omega_{Q}$ induced by the inner product ${\langle x,y\rangle}_{Q}=x^{\top}Qy$.
  • Last, we may consider the elements $\Omega$ of the symplectic matrix group [30] ${\mathrm{Sp}}(2n)$:
    ${\mathrm{Sp}}(2n)=\left\{\Omega={\left[\begin{array}[]{ll}A&B\cr C&D\end{array}\right]}{\ :\ }-C^{\top}A+A^{\top}C=0,-C^{\top}B+A^{\top}D=I,-D^{\top}B+B^{\top}D=0\right\},$
    to define a corresponding symplectic form $\omega_{\Omega}$. $\Omega_{0}\in{\mathrm{Sp}}(2n)$ is the symplectic matrix corresponding to the canonical form $\omega_{0}$ of $\mathbb{R}^{2n}$. Transpose and inverse of symplectic matrices are symplectic matrices. Moreover, the inverse of a symplectic matrix $\Omega$ is $\Omega^{-1}=-\Omega_{0}\Omega^{\top}\Omega_{0}$. Symplectic matrices of ${\mathrm{Sp}}(2n)$ have unit determinant, and when $n=1$, ${\mathrm{Sp}}(2)$ corresponds precisely to the set of matrices with unit determinant.
    A transformation $t:V\rightarrow V$ is called a linear symplectomorphism when $\omega(t(z_{1}),t(z_{2})=\omega(z_{1},z_{2})$ (i.e., $t^{*}\omega=\omega$). Let $T$ be the matrix representation of $t$. Then $t$ is a linear symplectomorphism with respect to the canonical basis if $T^{\top}\Omega_{0}T=\Omega_{0}$.
3 Symplectic Fenchel transform, symplectic subdifferentials, and symplectic Fenchel-Young inequality
Let $F:Z=X\times Y\rightarrow\mathbb{R}\cup\{+\infty\}$ be a convex lower semicontinuous function called a potential function. The symplectic Fenchel conjugate $F^{*\omega}(z^{\prime})$ corresponding to $F(z)<+\infty$ is defined by
$F^{*\omega}(z^{\prime})=\sup_{z\in Z}\left\{\omega(z^{\prime},z)-F(z)\right\}.$
Remark 1
Moreau generalized the Fenchel conjugate using a cost function [24]. The duality induced by logarithmic cost function was studied in [33] and lead to a generalization of Bregman divergences called the logarithmic divergences.
The symplectic subdifferential of $F$ at $z$ is defined by
$\partial^{\omega}F(z)=\left\{z_{1}\in Z{\ :\ }\forall z_{2}\in Z,F(z+z_{2})\geq F(z)+\omega(z_{1},z_{2})\right\}.$
The differential operator $\partial^{\omega}$ is a set-valued operator: $\partial^{\omega}:{\mathcal{F}}\rightrightarrows Z$, where ${\mathcal{F}}$ is the set of potential functions. An element of the symplectic subdifferential of $F$ at $z$ is called a symplectic subgradient.
Remark 2
In geometric mechanics [9], the symplectic gradient on a symplectic manifold $(M,\omega)$ is the Hamiltonian vector field, i.e., the vector field $X_{H}$such that the Halmitonian mechanics write concisely as $\omega(X_{H},\cdot)=\mathrm{d}H$.
Theorem 1 (Symplectic Fenchel-Young inequality (Theorem 2.3 of [6]))
Let $F(z)$ be a convex and lower semicontinuous function. Then the following inequality holds
$\forall z,z^{\prime}\in Z,\quad F(z)+F^{*\omega}(z^{\prime})\geq\omega(z^{\prime},z),$
with equality if and only if $z^{\prime}\in\partial^{\omega}(z)$.
Let us assume that the potential functions are smooth and that symplectic differentials consist only of single-element sets. By abuse of language, we shall call in this paper the symplectic gradient of $F$ the single element of $\partial^{\omega}$ denoted by $\nabla^{\omega}F$ (thus not to be confuse with the Hamiltonian vector field $X_{H}$ of geometric mechanics).
4 Symplectic Fenchel-Young divergences and symplectic Bregman divergences
From the symplectic Fenchel-Young inequality, we can define the symplectic Fenchel-Young divergence as follows:
Definition 3 (Symplectic Fenchel-Young divergence)
$Y_{F}(z,z^{\prime})=F(z)+F^{*\omega}(z^{\prime})-\omega(z^{\prime},z)\geq 0.$ (6)
We have $Y_{F}(z,z^{\prime})=0$ if and only if $z^{\prime}\in\partial^{\omega}F(z)$.
Let us define the symplectic Bregman divergence $B_{F}^{\omega}(z_{1}:z_{2})$ as $Y_{F}(z,z^{\prime}_{2})$ where $z^{\prime}_{2}=\nabla^{\omega}F(z_{2})$. Using the identity $F(z)+F^{*\omega}(\nabla^{\omega}F(z))=\omega(\nabla^{\omega}F(z),z)$, we get
$B_{F}^{\omega}(z_{1}:z_{2})$ $=$ $Y_{F}(z,z^{\prime}_{2}),$
$=$ $F(z_{1})-F(z_{2})-\omega(\nabla^{\omega}F(z_{2}),z_{1}-z_{2}).$
Since $\omega$ is skew-symmetric, we can rewrite equivalently Eq. 4 as
$B_{F}^{\omega}(z_{1}:z_{2})=F(z_{1})-F(z_{2})+\omega(z_{1}-z_{2},\nabla^{\omega}F(z_{2})).$ (7)
Definition 4 (Symplectic Bregman divergence)
Let $(Z,\omega)$ be a symplectic vector space. Then the symplectic Bregman divergence between $z_{1}$ and $z_{2}$ of $Z$ is
$B_{F}^{\omega}(z_{1}:z_{2}):=F(z_{1})-F(z_{2})-\omega(\nabla^{\omega}F(z_{2}),z_{1}-z_{2}),\quad\partial^{\omega}F(z)=\{\nabla^{\omega}F(z)\}.$
Notice that ordinary Bregman divergences (BDs) have been generalized to non-smooth strictly convex potential functions using a subdifferential map in [17, 13, 15]. Similarly, we can extend symplectic BDs to non-smooth strictly convex potential functions.
5 Particular case: Even-dimensional composite Bregman divergences
When $Y=X$ and $(X,{\langle\cdot,\cdot\rangle})$ is an inner-product space, we may consider the composite inner-product on $Z$:
${\langle\langle z_{1},z_{2}\rangle\rangle}={\langle x_{1},y_{1}\rangle}+{\langle x_{2},y_{2}\rangle}.$
Let $I:Z\rightarrow Z$ be the linear function $I(z)=z$ and denote by $J:Z\rightarrow Z$ the linear function such that $J(z)=J(x,y)=(-y,x)$. Notice that this definition of $J$ makes sense because $X=Y$. We check that we have $J^{2}(x,y)=(-x,-y)$, i.e., $J^{2}=-I$. Furthermore, we have $g(z_{1},z_{2})=\omega(z_{1},Jz_{2})={\langle x_{1},x_{2}\rangle}+{\langle y_{1},y_{2}\rangle}$ that is a positive definite inner product. That is, the automorphism $J$ is a complex structure $\omega$-compatible ($J$ is a symplectomorphism).
We can express the symplectic form $\omega$ using the composite inner product as follows:
$\omega(z_{1},z_{2})$ $=$ ${\langle\langle J(z_{1}),z_{2}\rangle\rangle},$ (8)
$\omega(-Jz_{1},z_{2})$ $=$ ${\langle\langle z_{1},z_{2}\rangle\rangle}.$ (9)
Similarly, the symplectic subdifferential of $F$ can be expressed using the ordinary subdifferential as follows:
$z^{\prime}\in\partial^{\omega}F(z)$$\Leftrightarrow$ $J(z^{\prime})\in\partial F(z),$ (10)
$z^{\prime}\in\partial F(x)$$\Leftrightarrow$ $-J(z^{\prime})\in\partial^{\omega}F(z).$ (11)
Last, the symplectic Fenchel conjugate is related by the ordinary Fenchel conjugate:
$F^{*\omega}(z)=F^{*}(J(z)).$
Thus in that case the symplectic Bregman divergence amounts to an ordinary Bregman divergence:
$B^{\omega}(z_{1}:z_{2})=B_{F}(z_{1}:z_{2}).$
6 Summary, discussion, and perspectives
Since it inception in operations research, Bregman divergences [5] have proven instrumental in many scientific fields including information theory, statistics, and machine learning, just to cite a few. Let $(X,{\langle\cdot,\cdot\rangle})$ be a Hilbert space, and $F:X\rightarrow\mathbb{R}$ a strictly convex and smooth realvalue function. Then the Bregman divergence induced by $F$ is defined in [5] (1967) by
$B_{F}(x_{1}:x_{2})=F(x_{1})-F(x_{2})-{\langle x_{1}-x_{2},\nabla F(x_{2})\rangle}.$
In information geometry [1, 25, 26], a smooth dissimilarity $\mathcal{D}(p,q)$ between two points $p$ and $q$ on a $n$-dimensional smooth manifold $M$ induces a statistical structure on the manifold [11], i.e., a triplet $(g,\nabla,\nabla^{*})$ where the Riemannian metric tensor $g$ and the torsion-free affine connections $\nabla$ and $\nabla^{*}$ are induced by $\mathcal{D}$. The duality in information geometry is expressed by the fact that the mid-connection $\frac{\nabla+\nabla^{*}}{2}$ corresponds to the Levi-Civita connection induced by $g$. To build the divergence-based information geometry [2], the divergence $\mathcal{D}(p,q)$ is interpreted as a scalar function on the product manifold $M\times M$ of dimension $2n$. Thus the divergence $\mathcal{D}$ is called a contrast function [11] or yoke [4]. Conversely, a statistical structure $(g,\nabla,\nabla^{*})$ on a $n$-dimensional manifold $M$ induces a contrast function [21]. When the statistical manifold $(M,g,\nabla,\nabla^{*})$ is dually flat with $\theta(\cdot)$ the global $\nabla$-affine coordinate system and $\eta(\cdot)$ the global $\nabla^{*}$-affine coordinate system [29], there exists two dual potential functions $\phi$ and $\phi^{*}$ on $M$ such that $\phi(p)=F(\theta(p))$ and $\phi^{*}(\eta(p))=F^{*}(\eta(p))$ where $F^{*}(\eta)$ is the Legendre-Fenchel convex conjugate of $F(\theta)$. The canonical dually flat divergence on $M$ is then defined by
$\mathcal{D}(p,q)=F(\theta(p))+F^{*}(\eta(p))-\sum_{i=1}^{n}\theta_{i}(p)\eta_{i}(q),$
and amounts to a Fenchel-Young divergence or equivalently a Bregman divergence:
$\mathcal{D}(p,q)=Y_{F}(\theta(p):\theta(q))=B_{F}(\theta(p):\theta(q)),$
where the Fenchel-Young divergence is defined by
$Y_{F}(\theta:\eta^{\prime})=F(\theta)+F^{*}(\eta^{\prime})-\sum_{i=1}^{n}\theta^{i}\eta^{\prime}_{i}.$
It has been shown that a divergence $\mathcal{D}$ also allows one to define a symplectic structure on a statistical manifold [4, 34].
In this expository paper, we have defined symplectic Fenchel-Young divergences and equivalent symplectic Bregman divergences by following the study of geometric mechanics reported in [6]. The symplectic Bregman divergence between two points $z_{1}$ and $z_{2}$ on a symplectic vector space $(Z,\omega)$ induced by a convex potential function $F:Z\rightarrow\mathbb{R}$ is defined by
$B_{F}^{\omega}(z_{1}:z_{2})=F(z_{1})-F(z_{2})+\omega(z_{1}-z_{2},\nabla^{\omega}F(z_{2})),$
where $\nabla^{\omega}F$ has been called the symplectic gradient in this paper, and assumed to be the unique symplectic subdifferential at any $z\in Z$, i.e., $\partial^{w}F(z)=\{\nabla^{\omega}F(z)\}$.
In finite $2n$-dimensional symplectic vector spaces, linear symplectic forms $\omega$ can be represented by symplectic matrices of the matrix group ${\mathrm{Sp}}(2n)$. Buliga and de Saxcé [6] considered geometric mechanics with dissipative terms, and stated the following “Brezis-Ekeland-Nayroles principle” (BEN principle for short):
Definition 5 (BEN principle [6])
An evolution curve $z(t)=z_{\mathrm{rev}}(t)+z_{\mathrm{irr}}(t)\in Z$ in a geometric mechanic system with convex dissipation potential $\phi(z)$ satisfies $Y_{F}^{\omega}(\dot{z}(t),\dot{z}_{\mathrm{irr}}(t))=0$, where $Y_{F}^{\omega}$ is the symplectic Fenchel-Young divergence induced by $\phi$, and $z_{\mathrm{rev}}(t)$ and $z_{\mathrm{irr}}(t)$ are the reversible and irreversible parts of the particle $z(t)$.
The decomposition of $z=z_{\mathrm{rev}}+z_{\mathrm{irr}}$ into two parts can be interpreted as the Moreau’s proximation [23, 27] associated to the potential function $\phi$: Indeed, let $F(z)$ be a convex function of $Z=\mathbb{R}^{d}$. Then for all $z\in\mathbb{R}^{d}$, we can uniquely decompose $z$ as $z=z+z^{*}$ such that $F(z)+F^{*}(z^{*})={\langle z,z*\rangle}$ (Fenchel-Young equality) where $z^{*}=\nabla F(z)$ (see Proposition in §4 of [23]). The part $z$ is called the proximation with respect to $F$, and the part $z^{*}$ is the proximation with respect to the convex conjugate $F^{*}$.
We may consider the non-separable potential functions $F(z)=F(x,y)=xf(y/x)$ which are obtained from the perspective transform [10] of arbitrary convex functions $f(u)$ to define symplectic Bregman divergences. The perspective functions $F(x,y)$ are jointly convex if and only if their corresponding generators $f$ are convex. Such perspective transforms play a fundamental role in information theory [8] and information geometry [1].
In machine learning, symplectic geometry has been used for designing accelerated optimization methods [16, 28] (Bregman-Lagrangian framework) and physics-informed neural networks [20, 7] (PINNs).
This paper aims to spur interest in either designing or defining symplectic divergences from first principles, and to demonstrate their roles when studying the learning dynamics of ML and AI systems.